# Unsupervised learning in particle physics

### v1

@@ 3.2 Variational autoencoders
The first term is the reconstruction loss or the expected negative log-likelihood for the $i$-th data point, given the latent variable $z$. The expectation is taken with respect to the encoder’s distribution over the representations. The second term is the Kullback-Leibler divergence (see Sec. Appendix A Kullback-Liebler divergence for more details) between the encoder’s distribution $\pi_{\theta}(z|x)$ and the prior $\rho(z)$:
@@
In Ref. [42], the idea of using AEs and VAEs for online selection in the L1T is explored. L1T mainly uses theory-motivated simple requirements like the minimum energies of reconstructed leptons, jets, etc. The authors propose that an anomaly detection algorithm deployed in the LIT will be more effective as it will collect data in a model-independent way. The recent development of the hls4ml open-source library has allowed the authors to translate the AD algorithms on the FPGAs firmware mounted on the L1T boards. They use a DNN AE and a CNN AE separately on FPGAs. Ref. [42] instead of using the standard loss function as the anomaly score, use the $\mathbb{KL}$ divergence term and the $Z$-score of the origin $\vec{0}$ in the latent space with respect to the Gaussian distribution with standard deviation $\vec{\sigma}$ and centred at $\vec{\mu}$($\vec{\sigma}$ and $\vec{\mu}$ values are returned by the encoder). This helps in saving resources and latency by not having to evaluate the decoder in the trigger. There is also no need to do Gaussian sampling on the hardware, which will preserve the deterministic output of the trigger. They demonstrate that BSM signatures can be enhanced by three orders of magnitude while staying within the strict latency and resource constraints of a typical LHC event filtering system. The anomaly detection is shown to be performed in as little as $80$ ns using less than $3$% of the logic resources in the Xilinx Virtex VU9P FPGA. This capability to perform real-time anomaly detection on FPGAs has generated attention within the community and has opened the way to use this idea during the next data-taking campaign at the LHC.
The CMS experiment has recently used an ML-based trigger algorithm, AXOL1TL (using a VAE, dense neural network) [43], which selects anomalous events in real time. It was trained with unbiased data collected by the CMS experiment in 2023 at a centre-of-mass energy of $\sqrt{s}$=13.6 TeV and was implemented in Level-1 Global Trigger Test Crate for 2023 data taking. The Global Trigger Test Crate is a copy of the main Global Trigger system—it receives the same input data but does not trigger the readout of CMS. This provides a platform for thoroughly testing new trigger algorithms on live data without interrupting data taking. The AXOL1TL algorithm has shown an improvement of the signal efficiency by up to $46$% for the BSM signal of a Higgs decaying to two (pseudo)scalars of mass $15$ GeV each of which decays to bottom quark pairs. Their results show that the Anomaly Detection trigger built on Xilinx Virtex-7 FPGA operates at an extremely low latency of $50$ ns and takes up only a small fraction of the resources.
@@
$L[f_{1},f_{2},g_{1},g_{2}]$ $=$ $\sum_{i}R_{1}(x_{i})^{2}+\sum_{i}R_{2}(x_{i})^{2}+\lambda\leavevmode\nobreak\ \mathrm{DisCo}^{2}[R_{1}(X),R_{2}(X)],$ (17)
@@
$L(\theta,\varphi)=-\mathbb{E}_{z\approx q_{\theta}(z|x_{i})}[\log p_{\phi}(x_{i}|z)]+\beta\leavevmode\nobreak\ \mathbb{KL}(q_{\theta}(z|x_{i})||p(z));$ (19)
@@ 3.3 Weak supervision
For an illustration, let us consider the example given in Ref. [50]. The benchmark signal considered there is the production of $3$-TeV $W^{\prime}$ bosons in the fully hadronic channel. The heavy $W^{\prime}$ bosons decay to a $W$ boson and a new scalar $X$ of mass about $400$ GeV (i.e., $2m_{W}\ll M_{X}\ll M_{W^{\prime}}$), which further decays to a pair of $W$ bosons: $pp\to W^{\prime}\to WX,\;X\to W^{+}W^{-}$. The final-state signature is taken as two fatjets, where one of them is two-pronged ($W$ boson) and the other is four-pronged (the new scalar $X$), and QCD-dijets events are considered as the background. The CWoLa hunting method is evaluated on two datasets, one without any signal events and the other a mixture of signal and QCD dijet events. The feature of interest is $m_{JJ}$, the invariant mass of the two-fatjet system. The auxiliary variable, which has to be uncorrelated with $m_{JJ}$, is created from the mass of the fatjet and substructure variables: for each fatjet $J_{i}$, $Y_{i}=\left(m_{J},\sqrt{\tau^{\beta=2}_{1}}/\tau^{\beta=1}_{1},\tau^{\beta=1}_{21},\tau^{\beta=1}_{32},\tau^{\beta=1}_{43},n_{trk}\right)$, where $n_{trk}$ is the number of tracks associated with the fatjet and $\tau_{\alpha\delta}$’s are the $N$-subjettiness ratios [51] for $\alpha$-prongness check. Thus, the auxiliary variable becomes the $12$-features long, $Y\equiv(Y_{1},Y_{2})$. The SR and the SBR for the $m_{JJ}\simeq 3$ TeV hypothesis are chosen as in the left panel of Fig. 7. As seen in the right panel of Fig. 7, a clear bump is visible in the case of the signal-injected sample for different classifier thresholds, whereas no such peak is seen in the other mixture.
@@
Constructing Unobserved Regions by Transforming Adjacent Intervals (CURTAINs): In another technique called CURTAINs [56]—short for Constructing Unobserved Regions by Transforming Adjacent Intervals—a conditional invertible neural network (cINN, example can be seen in Fig. 14) is trained with an optimal transport loss to learn the background from the SBR. The cINNs are trained to transform data from the lower sideband of $m_{res}$ ($m_{JJ}$ in Ref. [56]) to the higher sideband in the forward pass and in the opposite direction in the inverse pass. These are conditioned on $m^{low}_{JJ}$ and $m^{high}_{JJ}$ (belonging to the respective sidebands) using the conditioning function $f(m^{low}_{JJ},m^{high}_{JJ})=m^{high}_{JJ}-m^{low}_{JJ}$. In the forward pass, $m^{low}_{JJ}$ is the true value of the input and $m^{high}_{JJ}$ becomes the target value. The authors pick the same dataset as described in the CWoLa hunting example before; the target values are sampled from a fit to QCD dijet background invariant mass distribution. The loss function is taken as the Sinkhorn divergence, which measures the distance between the original and the transformed distributions. After every epoch of training, an intra-sideband training is performed where each of the sidebands is divided into two halves and the network is trained to make transformations between those. The authors show that CURTAINs performs similarly to CATHODE in anomaly detection tasks using only the sideband region information and much less training data. A Normalising Flow-based update to CURTAINs can be found in Ref. [57].
Some hybrid techniques which makes of input from simulations to aid the unlabeled classification task. Simulation assisted likelihood-free anomaly detection (or SALAD for short) [58] uses a reweighting function to transform simulation into observed data parametrised on the auxilliary features. First, a classifier is trained to distinguish between simulated and the observed data in the SB. This classifier is used to buid a weighting function, $w(Y|m)$, which reweights the simulated events in the SR where $w$ is interpolated by a neural network into the SR. A second classifier is trained to distinguish the reweighted simulation from unlabeled data. The second classifier is monotonically related to the optimal classifier, and events can be selected by putting a high threshold on this classifier output. Another method, Flow-enhanced transportation for anomaly detection (FETA) [59] used normalising flows to transform simulation to observed data in the sideband regions. This transformation is used to create a simulation-informed background template in the SR.
Figure 10: Non-resonant anomaly detection setup from Ref. [60] showing signal and control regions. Background distribution learned in the CR needs to extroplated into the SR, unlike in resonant anomaly detection.
@@ 3.4 Topic modelling
As we saw in Section 3.3, CWoLa-based taggers trained on such mixtures could separate quark jets from gluon jets well, but they do not give any information about the underlying distributions $\rho_{k}(x)$. One can match these topic distributions $\rho_{\mathcal{T}_{i}}(x)$ to the actual underlying quark or gluon distributions, $\rho_{g}(x)$ and $\rho_{q}(x)$, if certain requirements are fulfilled by the samples. The main criterion is that the mixtures should have different purities, i.e., $f^{(a)}_{q}(x)$ for each mixture should be different. Another condition is the presence of anchor bins or mutual irreducibility of the distributions—an underlying distribution $\rho_{k}(x)$ cannot be a mixture of the rest of the distributions plus some other distribution. To illustrate the effectiveness of the algorithm, Ref. [62] generates the mixtures using some selection cuts: central jets ($|\eta|<2$) are selected, one for the $Z+\text{jets}$ events and two for the QCD dijet events if they have transverse momentum $p_{T}\in\left[250,275\right]$ GeV. At the truth level (Pythia labelled), these mixtures have $f_{q}^{(1)}=0.88$ and $f_{q}^{(2)}=0.37$. Fig. 11(b) shows the constituent multiplicity of the jets in each of these mixtures along with the Pythia-labelled distributions. The learned jet topics show good agreement with the truth-level distributions.
@@
3.5 Self-supervised learning methods
@@
Figure 12: JetCLR [68]: Alignment and Uniformity in the learned embeddings on $\mathbb{S}^{n-1}$ hypersphere with a contrastive loss. Alignment enforces that positive pairs have similar embeddings and therefore occupy the same location in the hypersphere. Uniformity ensures that the learned embeddings occupy the entire hypersphere to maximize the information learned [69].
@@
DarkCLR [73] extends AnomalyCLR to identify anomalous semi-visible jets at LHC. Semi-visible jets show up in models that incorporate the Dark sector and generally contain fewer constituents compared to usual jets. Therefore, the anomalous augmentation in this case involves dropping constituents from the jets. The authors also use a normalized autoencoder (NAE) to calculate the anomaly score. The method results in better performance compared to AnomalyCLR for low small values of $\epsilon_{B}^{-1}$.
@@
4 Detector unfolding and simulation
There are two aspects to testing the prospects of a NP model at the LHC. First, one needs precise simulations [^8] For example, one can use MC tools like MadGraph5 [74, 75] and Pythia [76] to produce events and study them within a detector environment with Delphes [77] or Geant4 [78, 79], etc. to test the viability of the model. Full simulations of the particle showers in calorimeters can be computationally expensive, sometimes taking minutes per event even on a high-performance platform. As the LHC enters the high-luminosity mode, simulations will become even more computationally expensive and the need for alternative approaches for fast simulation will increase. The other is the inverse process of mapping the observed data to hard scatterings accurately. This process is called unfolding and generally, it is highly challenging as it has to take real detector effects such as acceptance, smearing, threshold effects, etc. into account. Some matrix-based methods have been developed to perform unfolding of 1D histograms [80, 81]. However, these methods scale exponentially with the joint unfolding of multiple variables making them practically impossible. They also require binning which can lead to further problems like statistical uncertainties and figuring out the (near-)optimal binning. Deep generative models, especially generative adversarial networks (GANs) and their variations, offer promising solutions to these issues. We present a brief overview of how they work below. (For detailed reviews on GANs, see Refs. [82, 83].)
@@ 4 Detector unfolding and simulation
The learning in this game resembles minimizing the Jensen-Shannon divergence (JSD)(see Sec. Appendix B Jensen-Shannon divergence for more details) between the data and the model distribution, and the game converges to its equilibrium if both neural networks can be updated directly in parameter space.
4.1 Unfolding detector effects
@@
In Ref. [85], the authors test two GAN frameworks for unfolding: a naive one and a fully conditional GAN. For the naive GAN, the generator input is the detector response (measurement) and is trained to output the parton distribution. The discriminator is trained to identify between the generated distribution and the true distribution. One can note that, in this scenario, there is no need for an event-by-event matched training pair between the detector response and the true distribution. Additionally, to improve the performance of the generator and provide extra training signal, the authors calculate the Maximum Mean Discrepancy (MMD) distance between a batch of generated samples and the true parton distributions. This naive GAN can unfold events effectively as long as the training and test datasets are statistically identical. In general, one cannot assume that the training and test datasets are arbitrarily similar, so to resolve this, the authors introduce a fully conditional GAN (FCGAN). The idea behind a conditional setup is not to learn a deterministic link between the input and the output (as is the case in supervised learning), but rather to learn to sample from a structured distribution $\rho(x|y)$. In FCGAN, the training is modified such that the generator is provided a random vector $r$ along with the detector response as inputs to generate the parton distribution. The discriminator is also modified such that it also has access to the detector response when classifying between real and generated samples. In this case one does require an event-to-event matched training pairs of detector response and parton distribution. The authors do not modify the MMD loss to be conditional since there is no efficient implementation of the same. Fig. 13 illustrates the model setup and the training losses applied. In general, FCGAN outperforms the naive GAN, even at cases where the training and the test datasets are not identical. However, it fails to reconstruct the invariant mass distribution, which the authors hypothesise is due to the lack of a fully conditional MMD loss. Using FCGAN, the authors also test the applicability of a GAN trained on the SM for unfolding BSM events. They find that, usually, the model is able to reconstruct the parton distributions for the BSM process, though not as well as in the previous cases.
@@
In Ref. [86], authors utilize a conditional invertible neural network (cINN) to model the transformation from the detector response to the parton distribution and vice versa. The motivation for using an INN is that the extra training signal provided by testing the parton to detector simulation would improve performance. Invertible neural networks (INNs) are a class of neural networks that are designed such that the learned functions are invertible. They allow for information-preserving transformations and are effective at tackling probabilistic modelling and generative tasks. Specifically, unlike traditional neural networks that tackle ambiguous inverse problems, INNs focus on learning the forward process while implicitly capturing the reverse process. The authors first test a naive implementation of INNs using a MSE loss to check for the reconstruction on both sides and the MMD loss to improve reconstruction. Similar to the naive GAN, the naive INN fails to capture the hard cuts imposed by the jet definition, especially in the tails of the distributions. To mitigate these, the authors introduce a conditional INN (cINN) that works by linking the parton-level data ($x_{p}$) to random noise ($r$) conditioned on the detector-level data ($x_{d}$). Trained on a given process, this network can generate the particle-level distribution given the detector-level events and an unfolding model. While this two-way training does not represent the inversion of a detector, it stabilises the training by requiring the input/reconstructed noise to be Gaussian. A small subnet first processes the detector-level data to provide the conditioning, i.e., $x_{d}\rightarrow f(x_{d})$. This subnet is trained alongside the cINN and does not need to be reversed or adapted. Fig. 14 illustrates the conditional INN setup used and the losses applied. The cINN model outperforms the simple INN model in representing hard cuts on the tails of the input distribution. Furthermore, the cINN model appears to be more calibrated than FCGAN. This fully conditional setup outperforms both the naive INN setup as well as a noise-extended INN setup for all the test cases. Finally, the authors test the performance of cINN on processes with initial-state radiations which changes the number of final jets at the detector level. Despite the variable number of jets (from $2$ to $4$), the model is effectively able to unfold the quantities with excellent performance individually for events with 2 jets, 3 jets, and 4 jets.
@@
Other kinds of methods, such as neural empirical Bayes [93], have been developed that take advantage of developments in neural density estimators to perform unfolding by learning $\rho(x|y)$ and $\rho(x)$ to maximise $\rho(y)$ using the Bayes theorem. The learned likelihoods can be used to reweigh the simulations to obtain parton-level distributions. Another method mitigates the problem of unfolding by learning optimized observables that can be calculated directly from the detector measurements [94], and can be unfolded easily using traditional unfolding strategies.
4.2 Detector simulation
Generative models like GANs have been explored for detector simulations due to their ability to learn and replicate complex data distributions [95, 96]. In this section, we look at some prominent applications of these models in detail.

### v3

@@ III.2 Variational autoencoders
The first term is the reconstruction loss or the expected negative log-likelihood for the $i$-th data point, given the latent variable $z$. The expectation is taken with respect to the encoder’s distribution over the representations. The second term is the Kullback-Leibler divergence (see Appendix Appendix A Kullback-Liebler divergence for more details) between the encoder’s distribution $\pi_{\theta}(z|x)$ and the prior $\rho(z)$:
@@
In Ref. [45], the idea of using AEs and VAEs for online selection in the L1T is explored. L1T mainly uses theory-motivated simple requirements like the minimum energies of reconstructed leptons, jets, etc. The authors propose that an anomaly detection algorithm deployed in the LIT will be more effective as it will collect data in a model-independent way. The recent development of the hls4ml open-source library has allowed the authors to translate the AD algorithms on the FPGAs firmware mounted on the L1T boards. They use a DNN AE and a CNN AE separately on FPGAs. Instead of using the standard loss function as the anomaly score, Ref. [45] uses the $\mathbb{KL}$ divergence term and the $Z$-score of the origin $\vec{0}$ in the latent space with respect to the Gaussian distribution with standard deviation $\vec{\sigma}$ and centred at $\vec{\mu}$ ($\vec{\sigma}$ and $\vec{\mu}$ values are returned by the encoder). This helps in saving resources and latency by not having to evaluate the decoder in the trigger. There is also no need to do Gaussian sampling on the hardware, which will preserve the deterministic output of the trigger. They demonstrate that BSM signatures can be enhanced by three orders of magnitude while staying within the strict latency and resource constraints of a typical LHC event filtering system. The anomaly detection is shown to be performed in as little as $80$ ns using less than $3$% of the logic resources in the Xilinx Virtex VU9P FPGA. This capability to perform real-time anomaly detection on FPGAs has generated attention within the community and has opened the way to use this idea during the next data-taking campaign at the LHC.
The CMS experiment has recently used an ML-based trigger algorithm, AXOL1TL (using a VAE, dense neural network) [46], which selects anomalous events in real-time. It was trained with unbiased data collected by the CMS experiment in 2023 at a centre-of-mass energy of $\sqrt{s}$=13.6 TeV and was implemented in Level-1 Global Trigger Test Crate for 2023 data taking. The Global Trigger Test Crate is a copy of the main Global Trigger system—it receives the same input data but does not trigger the readout of CMS. This provides a platform for thoroughly testing new trigger algorithms on live data without interrupting data taking. The AXOL1TL algorithm has shown an improvement of the signal efficiency by up to $46$% for the BSM signal of a Higgs decaying to two (pseudo)scalars of mass $15$ GeV each of which decays to bottom quark pairs. Their results show that the Anomaly Detection trigger built on Xilinx Virtex-7 FPGA operates at an extremely low latency of $50$ ns and takes up only a small fraction of the resources.
@@
$L[f_{1},f_{2},g_{1},g_{2}]$ $=$ $\sum_{i}R_{1}(x_{i})^{2}+\sum_{i}R_{2}(x_{i})^{2}+\lambda~{}\mathrm{DisCo}^{2}[R_{1}(X),R_{2}(X)],$ (17)
@@
$L(\theta,\varphi)=-\mathbb{E}_{z\approx q_{\theta}(z|x_{i})}[\log p_{\phi}(x_{i}|z)]+\beta~{}\mathbb{KL}(q_{\theta}(z|x_{i})||p(z));$ (19)
@@ III.3 Weak supervision
For an illustration, let us consider the example given in Ref. [53]. The benchmark signal considered there is the production of $3$-TeV $W^{\prime}$ bosons in the fully hadronic channel. The heavy $W^{\prime}$ bosons decay to a $W$ boson and a new scalar $X$ of mass about $400$ GeV, which further decays to a pair of $W$ bosons (since $2m_{W}\ll M_{X}\ll M_{W^{\prime}}$): $pp\to W^{\prime}\to WX,\;X\to W^{+}W^{-}$. The final-state signature is taken as two fatjets, where one of them is two-pronged ($W$ boson) and the other is four-pronged (the new scalar $X$), and QCD-dijets events are considered as the background. The CWoLa hunting method is evaluated on two datasets, one without any signal events and the other a mixture of signal and QCD dijet events. The feature of interest is $m_{JJ}$, the invariant mass of the two-fatjet system. The auxiliary variable, which has to be uncorrelated with $m_{JJ}$, is created from the mass of the fatjet and substructure variables: for each fatjet $J_{i}$, $Y_{i}=\left(m_{J},\sqrt{\tau^{\beta=2}_{1}}/\tau^{\beta=1}_{1},\tau^{\beta=1}_{21},\tau^{\beta=1}_{32},\tau^{\beta=1}_{43},n_{trk}\right)$, where $n_{trk}$ is the number of tracks associated with the fatjet and $\tau_{\alpha\delta}$’s are the $N$-subjettiness ratios [54] for $\alpha$-prongness check. Thus, the auxiliary variable becomes the $12$-features long, $Y\equiv(Y_{1},Y_{2})$. The SR and the SBR for the $m_{JJ}\simeq 3$ TeV hypothesis are chosen as in the left panel of Fig. 6. As seen in the right panel of Fig. 6, a clear bump is visible in the case of the signal-injected sample for different classifier thresholds, whereas no such peak is seen in the other mixture.
@@
Constructing Unobserved Regions by Transforming Adjacent Intervals (CURTAINs): In another technique called CURTAINs [59]—short for Constructing Unobserved Regions by Transforming Adjacent Intervals—a conditional invertible neural network (cINN, an example can be seen in Fig. 13) is trained with an optimal transport loss to learn the background from the SBR. The cINNs are trained to transform data from the lower sideband of $m_{res}$ ($m_{JJ}$ in Ref. [59]) to the higher sideband in the forward pass and in the opposite direction in the inverse pass. These are conditioned on $m^{low}_{JJ}$ and $m^{high}_{JJ}$ (belonging to the respective sidebands) using the conditioning function $f(m^{low}_{JJ},m^{high}_{JJ})=m^{high}_{JJ}-m^{low}_{JJ}$. In the forward pass, $m^{low}_{JJ}$ is the true value of the input and $m^{high}_{JJ}$ becomes the target value. The authors pick the same dataset as described in the CWoLa hunting example before; the target values are sampled from a fit to QCD dijet background invariant mass distribution. The loss function is taken as the Sinkhorn divergence, which measures the distance between the original and the transformed distributions. After every epoch of training, intra-sideband training is performed where each of the sidebands is divided into two halves and the network is trained to make transformations between those. The authors show that CURTAINs method performs similarly to CATHODE in anomaly detection tasks using only the sideband region information and much less training data. A Normalising Flow-based update to CURTAINs can be found in Ref. [60].
Some hybrid techniques take input from simulations to aid the unlabeled classification task. Simulation-assisted likelihood-free anomaly detection (or SALAD for short) [61] uses a reweighting function to transform simulation into observed data parametrised on the auxiliary features. First, a classifier is trained to distinguish between simulated and observed data in the SB. This classifier is used to build a weighting function, $w(Y|m)$, which reweights the simulated events in the SR where $w$ is interpolated by a neural network into the SR. A second classifier is trained to distinguish the reweighted simulation from unlabeled data. The second classifier is monotonically related to the optimal classifier, and events can be selected by putting a high threshold on this classifier output. Another method, Flow-enhanced transportation for anomaly detection (FETA) [62] used normalising flows to transform simulation to observed data in the sideband regions. This transformation is used to create a simulation-informed background template in the SR.
Figure 9: Non-resonant anomaly detection setup from Ref. [63] showing signal and control regions. Unlike in resonant anomaly detection, the background distribution learned in the CR needs to be extrapolated into the SR.
@@ III.4 Topic modelling
As we saw in Section III.3, CWoLa-based taggers trained on such mixtures could separate quark jets from gluon jets well, but they do not give any information about the underlying distributions $\rho_{k}(x)$. One can match these topic distributions $\rho_{\mathcal{T}_{i}}(x)$ to the actual underlying quark or gluon distributions, $\rho_{g}(x)$ and $\rho_{q}(x)$, if certain requirements are fulfilled by the samples. The main criterion is that the mixtures should have different purities, i.e., $f^{(a)}_{q}(x)$ for each mixture should be different. Another condition is the presence of anchor bins or mutual irreducibility of the distributions—an underlying distribution $\rho_{k}(x)$ cannot be a mixture of the rest of the distributions plus some other distribution. To illustrate the effectiveness of the algorithm, Ref. [65] generates the mixtures using some selection cuts: central jets ($|\eta|<2$) are selected, one for the $Z+\text{jets}$ events and two for the QCD dijet events if they have transverse momentum $p_{T}\in\left[250,275\right]$ GeV. At the truth level (Pythia labelled), these mixtures have $f_{q}^{(1)}=0.88$ and $f_{q}^{(2)}=0.37$. Fig. 10(b) shows the constituent multiplicity of the jets in each of these mixtures along with the Pythia-labelled distributions. The learned jet topics show good agreement with the truth-level distributions.
@@
III.5 Self-supervised learning methods
@@
Figure 11: JetCLR [71]: Alignment and Uniformity in the learned embeddings on $\mathbb{S}^{n-1}$ hypersphere with a contrastive loss. Alignment enforces that positive pairs have similar embeddings and, therefore, occupy the same location in the hypersphere. Uniformity ensures that the learned embeddings occupy the entire hypersphere to maximize the information learned [72].
@@
DarkCLR [76] extends AnomalyCLR to identify anomalous semi-visible jets [77, 78, 79, 80, 81, 82, 83] at LHC. Semi-visible jets show up in models that incorporate the dark sector (from the general class of Hiddden Valley Models [33, 84, 85]) and generally contain fewer constituents compared to usual jets. Therefore, the anomalous augmentation, in this case, involves dropping constituents from the jets. The authors also use a normalized autoencoder (NAE) to calculate the anomaly score. The method results in better performance compared to AnomalyCLR for low small values of $\epsilon_{B}^{-1}$.
@@
IV Detector unfolding and simulation
There are two aspects to testing the prospects of a NP model at the LHC. First, one needs precise simulations [^8] For example, one can use MC tools like MadGraph5 [86, 87] and Pythia [88] to produce events and study them within a detector environment with Delphes [89] or Geant4 [90, 91], etc. to test the viability of the model. Full simulations of the particle showers in calorimeters can be computationally expensive, sometimes taking minutes per event even on a high-performance platform. As the LHC enters the high-luminosity mode, simulations will become even more computationally expensive and the need for alternative approaches for fast simulation will increase. The other is the inverse process of mapping the observed data to hard scatterings accurately. This process is called unfolding, and generally, it is highly challenging as it has to take real detector effects, such as acceptance, smearing, threshold effects, etc., into account. Some matrix-based methods have been developed to perform unfolding of 1D histograms [92, 93]. However, these methods scale exponentially with the joint unfolding of multiple variables making them practically impossible. They also require binning which can lead to further problems like statistical uncertainties and figuring out the (near-)optimal binning. Deep generative models, especially generative adversarial networks (GANs) and their variations, offer promising solutions to these issues. We present a brief overview of how they work below. (For detailed reviews on GANs, see Refs. [94, 95].)
@@ IV Detector unfolding and simulation
The learning in this game resembles minimizing the Jensen-Shannon divergence (JSD)(see Appendix Appendix B Jensen-Shannon divergence for more details) between the data and the model distribution, and the game converges to its equilibrium if both neural networks can be updated directly in parameter space.
IV.1 Unfolding detector effects
@@
A follow-up work on unfolding done in Ref. [97] also trains a GAN to perform unfolding. They test between two GAN frameworks: a naive one and a fully conditional GAN. For the naive GAN, the generator input is the detector response (measurement) and is trained to output the parton distribution. The discriminator is trained to identify between the generated distribution and the true distribution. One can note that, in this scenario, there is no need for an event-by-event matched training pair between the detector response and the true distribution. Additionally, to improve the performance of the generator and provide extra training signal, the authors calculate the Maximum Mean Discrepancy (MMD) distance between a batch of generated samples and the true parton distributions. This naive GAN can unfold events effectively as long as the training and test datasets are statistically identical. In general, one cannot assume that the training and test datasets are arbitrarily similar, so to resolve this, the authors introduce a fully conditional GAN (FCGAN). The idea behind a conditional setup is not to learn a deterministic link between the input and the output (as is the case in supervised learning), but rather to learn to sample from a structured distribution $\rho(x|y)$. In FCGAN, the training is modified such that the generator is provided with a random vector $r$ along with the detector response as inputs to generate the parton distribution. The discriminator is also modified such that it also has access to the detector response when classifying between real and generated samples. In this case one does require an event-to-event matched training pairs of detector response and parton distribution. The authors do not modify the MMD loss to be conditional since there is no efficient implementation of the same. Fig. 12 illustrates the model setup and the training losses applied. In general, FCGAN outperforms the naive GAN, even in cases where the training and the test datasets are not identical. However, it fails to reconstruct the invariant mass distribution, which the authors hypothesise is due to the lack of a fully conditional MMD loss. Using FCGAN, the authors also test the applicability of a GAN trained on the SM for unfolding BSM events. They find that, usually, the model is able to reconstruct the parton distributions for the BSM process, though not as well as in the previous cases.
@@
In a follow-up work [98], authors instead utilize a cINN to model the transformation from the detector response to the parton distribution and vice versa. Invertible neural networks (INNs) are a class of neural networks that are designed such that the learned functions are invertible. They allow for information-preserving transformations and are effective at tackling probabilistic modelling and generative tasks. Specifically, unlike traditional neural networks that tackle ambiguous inverse problems, INNs focus on learning the forward process while implicitly capturing the reverse process. The motivation for using an INN is that the extra training signal provided by testing the parton to detector simulation would improve performance. The authors first test a naive implementation of INNs using a MSE loss to check for the reconstruction on both sides and the MMD loss to improve reconstruction. Similar to the naive GAN, the naive INN fails to capture the hard cuts imposed by the jet definition, especially in the tails of the distributions. To mitigate these, the authors introduce a conditional INN (cINN) that works by linking the parton-level data ($x_{p}$) to random noise ($r$) conditioned on the detector-level data ($x_{d}$). Trained on a given process, this network can generate the particle-level distribution given the detector-level events and an unfolding model. While this two-way training does not represent the inversion of a detector, it stabilises the training by requiring the input/reconstructed noise to be Gaussian. A small subnet first processes the detector-level data to provide the conditioning, i.e., $x_{d}\rightarrow f(x_{d})$. This subnet is trained alongside the cINN and does not need to be reversed or adapted. Fig. 13 illustrates the conditional INN setup used and the losses applied. The cINN model outperforms the simple INN model in representing hard cuts on the tails of the input distribution. Furthermore, the cINN model appears to be more calibrated than FCGAN. This fully conditional setup outperforms both the naive INN setup as well as a noise-extended INN setup for all the test cases. Finally, the authors test the performance of cINN on processes with initial-state radiations which changes the number of final jets at the detector level. Despite the variable number of jets (from $2$ to $4$), the model is effectively able to unfold the quantities with excellent performance individually for events with 2 jets, 3 jets, and 4 jets.
@@
Other kinds of methods, such as neural empirical Bayes [105], have been developed that take advantage of developments in neural density estimators to perform unfolding by learning $\rho(x|y)$ and $\rho(x)$ to maximise $\rho(y)$ using the Bayes theorem. The learned likelihoods can be used to reweigh the simulations to obtain parton-level distributions. Another method mitigates the problem of unfolding by learning optimized observables that can be calculated directly from the detector measurements [106] and can be unfolded easily using traditional unfolding strategies.
IV.2 Detector simulation
Generative models like GANs have been explored for detector simulations due to their ability to learn and replicate complex data distributions [107, 108, 109, 110]. In this section, we look at some prominent applications of these models in detail.